OpenStudy (anonymous):

I need help. I do NOT understand this Algebra 2 college level stuff.

OpenStudy (anonymous):

okay what is the question?

OpenStudy (anonymous):

the question is Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (anonymous):

@terenzreignz @agent0smith @ganeshie8 @Hero @Callisto please help him

OpenStudy (anonymous):

im a girl

OpenStudy (anonymous):

sorry @Ab590 please help her

OpenStudy (anonymous):

its okay

OpenStudy (agent0smith):

The function f(x) is a difference of squares.... eg\[\Large f(x) = x^2 - 4\]

OpenStudy (agent0smith):

Can you factor it?

OpenStudy (anonymous):

I think so...

OpenStudy (agent0smith):

Use this\[\Large a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

yes so it would be \[(x+2) (x-2)\]

OpenStudy (agent0smith):

Correct. Now try the next one: The function g(x) is a sum of squares. a sum of squares is the same thing as a diff of squares, but with a plus sign not a minus sign between the two terms

OpenStudy (anonymous):

it would be like